Main INDEX
Monthly INDEX
PREV
NEXT
User name averett
Log entry time 21:08:05 on August 2,2001
Entry number 68663
keyword=daily summary
Summary of today's activities in preparation for beam:
Surveyed both spectrometers at 18.6 degrees
Set spectrometer momenta at 1.123 GeV (quasi-elastic)
Surveyed target ladder with oven on, then off, then on again.
Found that the target ladder shifts significantly when the oven is hot. This shift is enough to crash the target ladder into the pickup coils. The target position was adjusted with the oven warm and it will now clear the pickup coils. However, do not move the target if the oven is not hot.
Based on the survey, we have changed the target mover settings to place the target in the correct spot for beam. Note however, that the reference cell position was left unchanged because it would travel past the upper limit. The misalignment in this position is roughly 2.7mm, and the beam may need to be additionally steered in this position.
To properly set the beam position for normal running, it should be at (x,y)=0,-1.2mm on the BeO which should roughly correspond to (x,y)=1.5mm,0 on the BPM's. There is a roughly 2mm tilt from the front to back on the target ladder with the upstream end about 2mm higher than the downstream. With the coordinates above, the beam should pass through the center of the target volume.
EPR diode was aligned. The wire to the EPR coil was found to be broken at the scattering chamber connector and was subsequently cut, fed through the bottom of the scattering chamber, and soldered directly together.
Work on the helicity electronics was completed and will be reported in another HALOG entry.
Cameras were aligned and target chamber was closed at around 8:00 pm.
Around 8:45 a sweep of the hall was begun and once completed, MCC estimates 1-2 hours before stable beam is delivered.